# New Job, New Gender? Measuring the Social Bias in Image Generation Models

### v1

@@ Abstract.
Image generation models can generate or edit images from a given text. Recent advancements in image generation technology, exemplified by DALL-E and Midjourney, have been groundbreaking. These advanced models, despite their impressive capabilities, are often trained on massive Internet datasets, making them susceptible to generating content that perpetuates social stereotypes and biases, which can lead to severe consequences. Prior research on assessing bias within image generation models suffers from several shortcomings, including limited accuracy, reliance on extensive human labor, and lack of comprehensive analysis. In this paper, we propose BiasPainter, a novel metamorphic testing framework that can accurately, automatically and comprehensively trigger social bias in image generation models. BiasPainter uses a diverse range of seed images of individuals and prompts the image generation models to edit these images using gender, race, and age-neutral queries. These queries span 62 professions, 39 activities, 57 types of objects, and 70 personality traits. The framework then compares the edited images to the original seed images, focusing on any changes related to gender, race, and age. BiasPainter adopts a testing oracle that these characteristics should not be modified when subjected to neutral prompts. Built upon this design, BiasPainter can trigger the social bias and evaluate the fairness of image generation models. To evaluate the effectiveness of BiasPainter, we use BiasPainter to test five widely-used commercial image generation software and models, such as stable diffusion and Midjourney. Experimental results show that 100% of the generated test cases can successfully trigger social bias in image generation models. According to our human evaluation, BiasPainter can achieve 90.8% accuracy on automatic bias detection, which is significantly higher than the results reported in previous work. Additionally, BiasPainter provides interesting and valuable insights from evaluating the fairness of image generation models, which is helpful for developers to improve the models in the future. All the code, data, and experimental results will be released to facilitate future research.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXsubmissionid: 786copyright: none
@@ 1. Introduction
Image generation models, which generate images from a given text, have recently drawn a lot of interest from academia and the industry. For example, Stable Diffusion (Rombach et al., 2021a), an open-sourced latent text-to-image diffusion model, has 60K stars on github  [^1] https://github.com/CompVis/stable-diffusion. And Midjourney, an AI image generation commercial software product launched on July 2022, has more than 15 million users (Dawood, 2023). These models are capable of producing high-quality images that depict a variety of concepts and styles when conditioned on the textual description and can significantly facilitate content creation and publication.
@@
Despite the extraordinary capability of generating various vivid images, image generation models are prone to generate content with social bias, stereotypes, and even hate. For example, previous work has found that text-to-image generation models tend to associate males with software engineers, females with housekeepers, and white people with attractive people (Bianchi et al., 2022). It is because these models have been trained on massive datasets of images and text scraped from the web, which is known to contain stereotyped, biased, and toxic contents (Hugo Touvron, 2023). The ramifications of such biased content are far-reaching, from reinforcing stereotypes to causing brand and reputation damage and even impacting individual well-being.
Figure 1. Examples of Biased Images Generated from Image Generation Models
@@
To mitigate the social bias and stereotypes in image generation models, an essential step is to trigger the biased content and evaluate the fairness of such models. Previous works have designed methods to evaluate the bias in image generation models. For example, (Cho et al., 2022) generates images from a set of words that should not be related to a specific gender or race (e.g., secretary, rich person). Then, with a pre-trained image-text alignment model named CLIP (Radford et al., 2021), the authors classify the generated images into gender and race categories. However, these works suffer from several drawbacks. First, their accuracy is relatively low. Detecting race and gender is not an easy task, considering the high diversity of generating style and content. According to their human evaluation results, the detecting method is not accurate (e.g., (Cho et al., 2022) only achieves 40% on race), leading to concern about the effectiveness and soundness of the method. Second, their scope is limited. Previous work (Bianchi et al., 2022) involves human annotation to evaluate the bias in generated images, aiming for an accurate evaluation. However, this manual method needs extensive human effort and is not scalable. Third, there is a lack of comprehensive evaluation across various demographic groups. According to a previous study (Bianchi et al., 2022), more than 90% of the images produced by image generation models are of white people, which implies the comprehensive evaluation of the bias in other groups, such as East Asian people and black people.
@@
In this paper, we design a novel metamorphic testing framework, BiasPainter, that can automatically, comprehensively and accurately measure the social bias in image generation models. In particular, BiasPainter operates by inputting both images and text (prompts) into these models and then analyzing how the models edit these images. Under this setting, we define a metamorphic relation that when prompted with gender, race, and age-neutral prompts, the gender, race, and age of the original image should not significantly alter after editing. Specifically, BiasPainter first collects photos of people across different races, genders, and ages as seed images. Then, it prompts the model to edit each seed image. The prompts are selected from a pre-defined comprehensive gender/racial/age-neutral prompt list covering professions, objects, personalities and activities. After that, BiasPainter measures the changes from the seed image to the generated image according to race, gender, and age. An ideal case is that race, gender, and age do not change significantly under the editing with a gender, racial, and age-neutral prompt. On the other hand, if a model is prone to change significantly and consistently (e.g., increasing the age of the person in the original image) under a specific prompt (e.g., ”a photo of a mean person”), BiasPainter detects a biased association(e.g., between elder and mean).
@@
To evaluate the effectiveness of BiasPainter, we conduct experiments on five widely-deployed image generation models and software products, stable-diffusion 1.5, stable-diffusion 2.1, stable-diffusion XL, Midjourney and InstructPix2Pix. To consider the stochastic nature of generative AI, we adopt three photos from each combination of gender, race, and age, ending up with 54 seed images, and we adopt 228 prompts to edit each seed image. For each image generation model, we generate 54*228=12312 images to evaluate the bias. The results show that 100% of the generated test cases from BiasPainter can successfully trigger social bias in image generation models. In addition, based on human evaluation, BiasPainter can achieve an accuracy of 90.8% in identifying the bias in images, which is significantly higher than the performance reported in the previous work (40% on race) (Cho et al., 2022). Furthermore, BiasPainter can offer valuable insights into the nature and extent of biases within these models and serves as a tool for evaluating bias mitigation strategies, aiding developers in improving model fairness.
@@
  • We design and implement BiasPainter, the first metamorphic testing framework for comprehensively measuring the social biases in image generation models.
  • We perform an extensive evaluation of BiasPainter on five widely deployed commercial conversation systems and research models. The results demonstrate that BiasPainter can effectively trigger a massive amount of biased behavior.
  • We release the dataset, the code of BiasPainter, and all experimental results, which can facilitate real-world fairness testing tasks and further follow-up research.
Content Warning: We apologize that this article presents examples of biased images to demonstrate the results of our method. For the mental health of participating researchers, we prompted a content warning in every stage of this work to the researchers and annotators, and told them that they were free to leave anytime during the study. After the study, we provided psychological counseling to relieve their mental stress.
@@ 2.1. Image Generation Models
Image generation models, also known as Text-to-Image Generative Models, aim to synthetic images given natural language descriptions. There is a long history of image generation. For example, Generative Adversarial Networks  (Goodfellow et al., 2014) and Variational Autoencoders  (van den Oord et al., 2017), are two famous models that have been shown excellent capabilities of understanding both natural languages and visual concepts and generating high-quality images. Recently, diffusion models, such as DALL-E  [^2] https://openai.com/research/dall-e, Imagen  [^3] https://imagen.research.google/ and Stable Diffusion (Rombach et al., 2021b), have gained a huge amount of attention due to their significant improvements in generating high-quality vivid images. Despite the aforementioned work’s aim to improve the quality of image generation, it still remains uncertain whether these generative models contain more complex social biases and stereotypes.
Most of the currently used image generation models and software provide two manners of generating images. The first is generating images based on natural language descriptions only. The second manner is adopting an image editing manner that enables the user to input an image and then edit the image based on natural language descriptions. The former manner has more freedom while the latter one is more controllable.
@@ 2.2. Social Bias
Bias in AI systems has been a known risk for decades (Bordia and Bowman, 2019). As one of the most notorious biases, social bias is the discrimination for, or against, a person or group, compared with others, in a way that is prejudicial or unfair (Webster et al., 2022). It remains a complicated problem that is difficult to counteract. To study the social bias in machine learning software, the definitions of bias and fairness play a crucial role. Researchers and practitioners have proposed and explored various fairness definitions (Chen et al., 2022). The most widely used definition is statistical parity which requires the probability of a favorable outcome to be the same among different demographic groups. Formally, an AI system has the following two elements (Chakraborty et al., 2021):
  • A class label is called a favorable label if it gives an advantage to the receiver.
  • An attribute that divides the whole population into different groups.
For example, in the case of job application datasets, ”receive the job offer” is the favorable label, and according to the ”gender” attribute, people can be categorized into different groups, like ”male” and ”female”. The fairness of the AI system is defined as the goal that based on the attribute, different groups will be treated similarly to receive the favorable label. If not, the AI system is biased. This definition of bias and fairness is widely adopted in classification and regression tasks, where the favorable labels can be clearly assigned and the probabilities can be easily measured. However, this setting cannot be easily adopted for image generation models, since labels and probability are hard to measure for such models.
As one of the most important applications of AI techniques trained on massive Internet data, image generation models can inevitably be biased. Since such systems are widely deployed in people’s daily life, biased content generated by these systems, especially those related to social bias, may cause severe consequences. Social biased content is not only uncomfortable for certain groups but also can lead to a bad social atmosphere and even aggravate social conflicts. As such, exposing and measuring the bias in image generation models is a critical task. However, the automatic bias evaluation framework for image generation software and models is rarely studied by previous works, which motivates this work.
@@
2.3. Metamorphic Testing
Metamorphic testing (Chen et al., 2020) is a testing technique that has been widely employed to address the oracle problem. The core idea of metamorphic testing is to detect violations of metamorphic relations (MRs) across multiple runs of the software under test. Specifically, MR describes the relationship between input-output pairs of software. Given a test case, metamorphic testing transforms it into a new test case via a pre-defined transformation rule and then checks whether the corresponding outputs of these test cases returned by the software exhibit the expected relationship.
Metamorphic testing has been adapted to validate Artificial Intelligence (AI) software over the past few years. These efforts aim to automatically report erroneous results returned by AI software via developing novel MRs. In particular, Chen et al. (Chen et al., 2008) investigated the use of metamorphic testing in bioinformatics applications. Xie et al. (Xie et al., 2011) defined eleven MRs to test k-Nearest Neighbors and Naive Bayes algorithms. Dwarakanath et al. (Dwarakanath et al., 2018) presented eight MRs to test SVM-based and ResNet-based image classifiers. Zhang et al. (Zhang et al., 2018) tested autonomous driving systems by applying GANs to produce driving scenes with various weather conditions and checking the consistency of the system outputs.
@@ 3. Approach And Implementation
In this section, we present BiasPainter, a metamorphic testing framework designed for measuring the social bias in image generation models. BiasPainter uses photos of different persons as seed images and adopts various prompts to let image generation models edit the seed images. The key insight is that the gender/race/age of the person in the photo should not be modified too much under the gender/race/age-neutral prompts. Otherwise, a spurious correlation between gender/race/age and other properties (e.g. career) exists in the model. Namely, a suspicious bias is detected. For example, if an image generation model tends to convert more female photos to male photos under the prompt ”a photo of a lawyer”, a bias about lawyers on gender is detected. Figure 2 depicts the framework of BiasPainter, which consists of five stages:
@@
    Neutral Prompt List Collection: collect and annotate different prompts from various topics.
@@
    Image Generation: For each seed image, input different prompts to image generation models for generating images.
@@
    Properties Assessment: Adopt methods to access the race, gender and age properties for the seed image and generated images.
@@
    Bias Detection: Detect the potential social bias with the metamorphic relation.
@@ 3.1. Seed Image Collection
The first step of BiasPainter is collecting diverse photos of different people across various races, genders and ages. BiasPainter can adopt photos of any people as seed images. To limit the scope of experiments, we only consider 3 kinds of races (white people, black people, and east asian people), two genders (male and female), and three ages (young adult, middle age, and elderly) to conduct experiments. Specifically, we utilize a public-available dataset, VGGFace2 (Cao et al., 2018), and select 3 photos from each combination of gender, race and age, ending up with 3*2*3*3=54 seed images. We show the examples in Figure 3.
@@ 3.2. Neutral Prompt List Collection
BiasPainter adopts different prompts, such as ”a photo of a lawyer”, to edit the seed images. To achieve this, BiasPainter first collects diverse and comprehensive word lists, on the following four perspectives: profession (e.g. scientist), personality (e.g. rude), objects (e.g. gun), and activities (e.g. crying). These words were collected from varied resources. The profession-related words were collected from the U.S. Bureau of Labor Statistics [^4] https://www.bls.gov/emp/tables/emp-by-detailed-occupation.htm, and the words related to personality, objects and activities were collected from related dictionaries [^5] https://onlineteachersuk.com/personality-adjectives-list/ [^6] https://www.oxfordlearnersdictionaries.com/external/pdf/wordlists/oxford-3000-5000/American_Oxford_5000.pdf [^7] https://www.vocabulary.com/lists/189583. Since BiasPainter only adopts race-, gender-, and age-neutral prompts to edit the seed images, we then manually filter out the words that are race-, gender-, and age-related words, such as actor/actress and waiter/waitress. Note that race-neutral means that people of any race can relate to the prompts, such as being a scientist, being rude, with a gun, and crying. Specifically, we recruited 10 annotators, providing them with prompts we previously collected, and asked them to measure the words’ relevance to race, gender, or age on a scale from 1 (Strongly Irrelevant) to 5 (Strongly Relevant). We filter out the words with average relevancy higger than 3. Finally, we adopt four templates to generate prompts according to the topic. For example, BiasPainter adopts a person who is [crying] for the activity topic and a person with a [book] for the object topic. Table 1 shows the details of the final prompt lists, consisting of 228 prompts.
@@
Topic Template Number Examples
activity a person who is [word] 39 crawling, crying, fighting, knitting, laughing
object a person with a/an [word] 57 book, cigar, cleaner, gun, mansion, soccer
personality a person who is [word] 70 arrogant, brave, mean, nasty, reliable, selfish
profession a person who is a/an [word] 62 CEO, scientist, bus driver, lawyer, librarian

### v2

@@ Abstract.
Image generation models can generate or edit images from a given text. Recent advancements in image generation technology, exemplified by DALL-E and Midjourney, have been groundbreaking. These advanced models, despite their impressive capabilities, are often trained on massive Internet datasets, making them susceptible to generating content that perpetuates social stereotypes and biases, which can lead to severe consequences. Prior research on assessing bias within image generation models suffers from several shortcomings, including limited accuracy, reliance on extensive human labor, and lack of comprehensive analysis. In this paper, we propose BiasPainter, a novel evaluation framework that can accurately, automatically and comprehensively trigger social bias in image generation models. BiasPainter uses a diverse range of seed images of individuals and prompts the image generation models to edit these images using gender, race, and age-neutral queries. These queries span 62 professions, 39 activities, 57 types of objects, and 70 personality traits. The framework then compares the edited images to the original seed images, focusing on the significant changes related to gender, race, and age. BiasPainter adopts a key insight that these characteristics should not be modified when subjected to neutral prompts. Built upon this design, BiasPainter can trigger the social bias and evaluate the fairness of image generation models. We use BiasPainter to evaluate six widely-used image generation models, such as stable diffusion and Midjourney. Experimental results show that BiasPainter can successfully trigger social bias in image generation models. According to our human evaluation, BiasPainter can achieve 90.8% accuracy on automatic bias detection, which is significantly higher than the results reported in previous work.
Image Generation Models, Social Bias, Model Evaluation
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681433isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer vision
@@ 1. Introduction
Image generation models, which generate images from a given text, have recently drawn lots of interest from academia and the industry. For example, Stable Diffusion (Rombach et al., 2021a), an open-sourced latent text-to-image diffusion model, has 65K stars on github  [^1] https://github.com/CompVis/stable-diffusion. And Midjourney, an AI image generation commercial software product launched on July 2022, has more than 15 million users (Dawood, 2023). These models are capable of producing high-quality images that depict a variety of concepts and styles when conditioned on the textual description and can significantly facilitate content creation and publication.
@@
Despite the extraordinary capability of generating various vivid images, image generation models are prone to generate content with social bias, stereotypes, and even hate. For example, Google’s image generator, the Gemini, had generated a large number of images that were biased and contrary to historical facts, causing the service to be taken offline on an emergency basis (Milmo and Hern, 2024). Previous work has also found that text-to-image generation models tend to associate males with software engineers, females with housekeepers, and white people with attractive people (Bianchi et al., 2022). It is because these models have been trained on massive datasets of images and text scraped from the web, which is known to contain stereotyped, biased, and toxic contents (Hugo Touvron, 2023). The ramifications of such biased content are far-reaching, from reinforcing stereotypes to causing brand and reputation damage and even impacting individual well-being.
Figure 1. Examples of Biased Generation Detected by BiasPainter.
@@
To mitigate the social bias and stereotypes in image generation models, an essential step is to trigger the biased content and evaluate the fairness of such models. Previous works have designed methods to evaluate the bias in image generation models. For example, (Cho et al., 2022) generates images from a set of words that should not be related to a specific gender or race (e.g., secretary, rich person). Then, with a pre-trained image-text alignment model named CLIP (Radford et al., 2021), the authors classify the generated images into gender and race categories. However, these works suffer from several drawbacks. First, their accuracy is relatively low. Detecting race and gender is not an easy task, considering the high diversity of generating style and content. According to their human evaluation results, the detecting method is not accurate (e.g., (Cho et al., 2022) only achieves 40% on race), leading to concern about the effectiveness and soundness of the method. Second, their scope is limited. Previous work (Bianchi et al., 2022) involves human annotation to evaluate the bias in generated images, aiming for an accurate evaluation. However, this manual method needs extensive human effort and is not scalable. Third, there is a lack of comprehensive evaluation across various demographic groups. According to a previous study (Cho et al., 2022), more than 90% of the images produced by image generation models are of white people, which implies the comprehensive evaluation of the bias in other groups, such as East Asian people and black people.
@@
In this paper, we design a novel evaluation framework, BiasPainter, that can automatically, accurately and comprehensively measure the social bias in image generation models. In particular, BiasPainter operates by inputting both photos of people and text (prompts) into these models and then analyzing how the models edit these photos. A high-level idea is when prompted with gender, race, and age-neutral prompts, the gender, race, and age of the human in the photo should not significantly alter after editing. Specifically, BiasPainter first collects photos of people across different races, genders, and ages as seed images. Then, it prompts the model to edit each seed image. The prompts are selected from a pre-defined comprehensive gender/racial/age-neutral prompt list covering professions, objects, personalities and activities. After that, BiasPainter measures the changes from the seed image to the generated image according to race, gender, and age. An ideal case is that race, gender, and age do not change significantly under the editing with a gender, racial, and age-neutral prompt. On the other hand, if a model is prone to change significantly and consistently (e.g., increasing the age of the person in the original image) under a specific prompt (e.g., ”a photo of a mean person”), BiasPainter detects a biased association(e.g., between elder and mean).
@@
To evaluate the effectiveness of BiasPainter, we conduct experiments on six widely deployed image generation models: stable-diffusion 1.5, stable-diffusion 2.1, stable-diffusion XL, Dall-E 2, Midjourney and InstructPix2Pix. We sample three photos from each combination of gender, race, and age, ending up with 54 seed images, and adopt 228 prompts to edit each seed image. For each image generation model, we generate 54*228=12312 images and use the (original image, generated image) pairs as test cases to evaluate the bias. The results show that BiasPainter can successfully trigger social bias in image generation models. In addition, based on human evaluation, BiasPainter can achieve an accuracy of 90.8% in identifying the bias in images, which is significantly higher than the performance reported in the previous work (40% on race) (Cho et al., 2022). Furthermore, BiasPainter can offer valuable insights into the nature and extent of biases within these models and serves as a tool for evaluating bias mitigation strategies, aiding developers in improving model fairness.
@@
  • We design and implement BiasPainter, A novel evaluation framework for comprehensively measuring the social biases in image generation models.
  • We perform an extensive evaluation of BiasPainter on six widely deployed commercial conversation systems and research models. The results demonstrate that BiasPainter can effectively trigger a massive amount of biased behavior.
  • We release the dataset, the code of BiasPainter, and all experimental results, which can facilitate real-world fairness evaluation tasks and further follow-up research.
Content Warning: We apologize that this article presents examples of biased images to demonstrate the results of our method.
@@ 2.1. Image Generation Models
Image generation models, also known as Text-to-Image Generative Models, aim to synthetic images given natural language descriptions. There is a long history of image generation. For example, Generative Adversarial Networks  (Goodfellow et al., 2014) and Variational Autoencoders  (van den Oord et al., 2017), are two famous models that have been shown excellent capabilities of understanding both natural languages and visual concepts and generating high-quality images. Recently, diffusion models, such as DALL-E  [^2] https://openai.com/research/dall-e, Imagen  [^3] https://imagen.research.google/ and Stable Diffusion (Rombach et al., 2021b), have gained a huge amount of attention due to their significant improvements in generating high-quality vivid images. Despite the aforementioned work’s aim to improve the quality of image generation, such generative models are reported to contain social biases and stereotypes (Bianchi et al., 2022).
Most of the currently used image generation models provide two manners of generating images. The first is generating images based on natural language descriptions only. The second manner is adopting an image editing manner that enables the user to input an image and then edit the image based on natural language descriptions. The former manner has more freedom while the latter one is more controllable.
@@ 2.2. Social Bias
Bias in AI models has been a known risk for decades (Bordia and Bowman, 2019). As one of the most notorious biases, social bias is the discrimination for, or against, a person or group, compared with others, in a way that is prejudicial or unfair (Webster et al., 2022). To study the social bias in the machine learning models, the definitions of bias and fairness play a crucial role. Researchers and practitioners have proposed and explored various fairness definitions (Chen et al., 2022). The most widely used definition is statistical parity which requires the probability of a favorable outcome to be the same among different demographic groups. For example, in the case of job application datasets, ”receive the job offer” is favorable, and according to the ”gender” attribute, people can be categorized into different groups, like ”male” and ”female”. If ”male” and ”female” are treated similarly to ”receive the job offer”, the AI model is fair.
This definition of bias and fairness is widely adopted in classification and regression tasks, where the favorable labels can be clearly assigned and the probabilities can be easily measured. However, this setting cannot be easily adopted for image generation models, since labels and probability are hard to measure for such models.
As one of the most important applications of AI techniques trained on massive Internet data, image generation models can inevitably be biased. Since such models are widely deployed in people’s daily lives, biased content generated by these systems, especially those related to social bias, may cause severe consequences. Social biased content is not only uncomfortable for certain groups but also can lead to a bad social atmosphere and even aggravate social conflicts. As such, exposing and measuring the bias in image generation models is a critical task.
@@ 3. Approach And Implementation
In this section, we present BiasPainter, our evaluation framework for measuring the social bias in image generation models. BiasPainter uses photos of different persons as seed images and adopts various prompts to let image generation models edit the seed images. The key insight is that the gender/race/age of the person in the photo should not be modified too much under the gender/race/age-neutral prompts. Otherwise, a spurious correlation between gender/race/age and other properties exists in the model. Namely, a suspicious bias is detected. For example, if an image generation model tends to convert more female photos to male photos under the prompt ”a photo of a lawyer”, a bias about lawyers on gender is detected. Figure 2 depicts the framework of BiasPainter, which consists of five stages:
@@
    Neutral Prompt List Collection: collect and annotate different prompts from various domains.
@@
    Image Generation: Edit each seed image with different prompts and get generated images.
@@
    Properties Assessment: Assess the race, gender and age properties for the seed image and generated images.
@@
    Bias Detection: Measure the property changes to detect the social bias.
@@ 3.1. Seed Image Collection
The first step of BiasPainter is collecting diverse photos of different people across various races, genders and ages. To limit the scope of experiments, we only consider 3 kinds of races (white people, black people, and East Asian people), two genders (male and female), and three ages (young adult, middle age, and elderly) to conduct experiments. Specifically, we utilize a public-available dataset, VGGFace2 (Cao et al., 2018), and select 3 photos from each combination of gender, race and age, ending up with 3*2*3*3=54 seed images. Each image was manually selected and inspected to ensure the race, gender and age attributes aligned with the people within the image. We show the examples in the supplementary material.
@@ 3.2. Neutral Prompt List Collection
BiasPainter adopts different prompts, such as ”a photo of a lawyer”, to edit the seed images. To achieve this, BiasPainter first collects diverse and comprehensive word lists, on the following four domains: profession (e.g. scientist), personality (e.g. rude), objects (e.g. gun), and activities (e.g. crying). These words were collected from varied resources. The profession-related words were collected from the U.S. Bureau of Labor Statistics [^4] https://www.bls.gov/emp/tables/emp-by-detailed-occupation.htm, and the words related to personality, objects and activities were collected from related dictionaries [^5] https://onlineteachersuk.com/personality-adjectives-list/ [^6] https://www.oxfordlearnersdictionaries.com/external/pdf/wordlists/oxford-3000-5000/American_Oxford_5000.pdf [^7] https://www.vocabulary.com/lists/189583. Since BiasPainter only adopts race-, gender-, and age-neutral prompts to edit the seed images, we then manually filter out the words that are race-, gender-, and age-related, such as actor/actress and waiter/waitress. Note that race-neutral means that people of any race can relate to the prompts, such as being a scientist, being rude, with a gun, and crying. Specifically, we recruited 10 annotators, providing them with prompts we previously collected, and asked them to measure the words’ relevance to race, gender, or age on a scale from 1 (Strongly Irrelevant) to 5 (Strongly Relevant). We filter out the words with average relevancy higger than 3. Finally, we adopt four templates to generate prompts according to the domain. For example, BiasPainter adopts a person who is [crying] for the activity domain and a person with a [book] for the object domain. Table 1 shows the details of the final prompt lists, consisting of 228 prompts.
@@
Domain Template Num. Examples
activity a person who is [word] 39 crying, fighting, knitting
object a person with a/an [word] 57 cigar, gun, soccer
personality a person who is [word] 70 arrogant, brave, mean
profession a person who is a/an [word] 62 CEO, scientist, lawyer